OpenStudy (anonymous):

PLEASE HELP Beatrice is considering a 3/27 balloon mortgage with an interest rate of 4.25% to purchase a house for $197,000. What will be her balloon payment at the end of 3 years? A) $186,598.50 B) $159,245.98 C) $150,118.49 D) $225,368.29 @studyshark2000 please? Or anyone

OpenStudy (amistre64):

the remaining balance ... ive got a formula that can determine it but previous interactions havent been that productive at times . \[B_n=Bk^n-P\frac{k^n-1}{k-1}\] we would have to use the formula to determine payments, and then we can use it to determine the balance after 3 years

OpenStudy (amistre64):

\[P=Bk^{27*12}\frac{k-1}{k^{27*12}-1}\] using this to find the 3*12th balance period \[B_{36}=Bk^{36}-Bk^{324}\frac{k-1}{k^{324}-1}\frac{k^{36}-1}{k-1}\] \[B_{36}=Bk^{36}(1-k^{288}\frac{k^{36}-1}{k^{324}-1})\]

OpenStudy (anonymous):

@amistre64 thankyou but i have no calculator on me to try and solve the equation and i wouldnt know where to start either but thankyou

OpenStudy (amistre64):

i know, how are you spose to work the solutions tho? you have to have some tools available to you for the course.

OpenStudy (amistre64):

tables even ... but those are a pain

OpenStudy (amistre64):

and you say wolfram doesnt open for you ?

OpenStudy (anonymous):

I know, youre right. Its just my sister needed it for geometry and i gave it to her. Forgot to ask for it back and i always forget to ask for it until i need it. && whats wolfram? @amistre64

OpenStudy (amistre64):

its an online math calculator, or at least thats what i use it for

OpenStudy (amistre64):

i use it to run the calculations with so i dont have to make so many erros imputing it into a calculator http://www.wolframalpha.com/input/?i=197000%281%2B.0425%2F12%29%5E%2836%29-1024%28%281%2B.0425%2F12%29%5E%2836%29-1%29+%2F+%281%2B.0425%2F12-1%29

OpenStudy (amistre64):

i was trying to think of a way to approximate it or at least narrow it down .. assuming no interest 197000/27, times 3 = 21889 197000 - 21889 is the least amount owed if there was not interest applied. so the solution has to be greater than 175111, but that just gets rid of 2 of them

OpenStudy (amistre64):

ive got no good reason for this next approach other than it gets us closer to the solution 4.25% is about 4% 175111(1.04) = 182115.44

OpenStudy (anonymous):

@amistre64 i knooow maybe i should just put $186,598.50?

OpenStudy (amistre64):

id be ok with that ..... no argument from me :)